The demand equation for the Drake GPS Navigator is
x + 4p − 1100 = 0,
where x is the quantity demanded per week and p is the wholesale unit price in dollars. The supply equation
x − 21p + 1000 = 0,
where x is the quantity the supplier will make available in the market each week when the wholesale price is p dollars each. Find the equilibrium quantity and the equilibrium price for the GPS Navigators.
equilibrium quantity= units
equilibrium price $=

Answer by ankor@dixie-net.com(22740)

The demand equation for the Drake GPS Navigator is: x + 4p - 1100 = 0,
where x is the quantity demanded per week and p is the wholesale unit price in dollars.
:
The supply equation
x - 21p + 1000 = 0,
where x is the quantity the supplier will make available in the market each week when the wholesale price is p dollars each.
:
Find the equilibrium quantity and the equilibrium price for the GPS Navigators.
equilibrium quantity= units
equilibrium price $=
:
Equilibrium occurs when demand = supply, therefore
x + 4p - 1100 = x - 21p + 1000
Subtract x from both sides
4p - 1100 = -21p + 1000
4p + 21p = 1000 + 1100
25p = 2100
p = 2100/25
p = $84 equilibrium prices
;
Find x using the demand equation
x + 4(84) - 1100 = 0
x + 336 - 1100 =
x - 764 = 0
x = 764 units equilibrium qty
:
:
Check solutions in the supply equation; x -21p + 1000 = 0
764 - 21(84) + 1000 = 0
764 - 1764 + 1000 = 0
-